Grade 9Magical Mathematics[Interesting Approach]

I Am In Class 9

Please Help Me With My Doubts

If x = 1+51/3+52/3 then find the value of x3-3x2-12x+6

PRAKHAR  GUPTA
13 Years agoGrade 9

3 Answers

Shivam Dimri
13 Years ago

its actually a geometric progression

x = (24/(51/3 - 1))

put it in the expression

.....

its too wierd i beilive!!

Akash Kumar Dutta
ApprovedApproved Tutor Answer13 Years ago

(x-1)^3=5+25 + 3.5(x-1)
then simplyfy

Abhishekh kumar sharma
13 Years ago

we know x=1+5^1\3+5^2\3

then we have x-1=a+a^2 where a=5^1\3

then put (x-1)^3=a+a^2

and solve